OpenStudy (anonymous):

The first four terms of a sequence are shown below: 7, 4, 1, -2 Which of the following functions best defines this sequence? f(1) = 7, f(n + 1) = f(n) + 3; for n ≥ 1 f(1) = 7, f(n + 1) = f(n) - 3; for n ≥ 1 f(1) = 7, f(n + 1) = f(n) - 4; for n ≥ 1 f(1) = 7, f(n + 1) = f(n) + 4; for n ≥ 1

OpenStudy (anonymous):

@texaschic101 plz help me

OpenStudy (anonymous):

guess: 2nd one cause it's number -3 to get next in sequence

OpenStudy (anonymous):

7-3=4 4-3=1 1-3=-2

OpenStudy (anonymous):

C?

OpenStudy (anonymous):

xq6